Suppose that a random sample of fifteen recently sold houses in a certain city has a mean sales price of $285,000, with a standard deviation of $5000. Under the assumption that house prices are normally distributed, find a 90% confidence interval for the mean sales price of all houses in this community.
Carry your intermediate computations to at least three decimal places. Round your answers to the nearest whole number
What is the lower limit of the confidence interval?
What is the upper limit of the confidence interval?
I think i knew how to do this once, but am not sure what formula to use now. Any help would be greatly appreciated!
Thank you in advance.

Answer by stanbon(75887)

You can find all this statistics stuff using Google.
-------------
CI: formula x-bar-E < u < x-bar+E
-----------------------------------
Suppose that a random sample of fifteen recently sold houses in a certain city has a mean sales price of $285,000, with a standard deviation of $5000. Under the assumption that house prices are normally distributed, find a 90% confidence interval for the mean sales price of all houses in this community.
Carry your intermediate computations to at least three decimal places. Round your answers to the nearest whole number
What is the lower limit of the confidence interval?
What is the upper limit of the confidence interval?
-----------
x-bar = 285000
E = t*s/sqrt(n)
E = 1.761*5000/sqrt(15) = 2273.44
285000-2273.44 < u < 285000+2273.44
==================================
Cheers,
Stan H.